OpenStudy (anonymous):

help me on x=11y+4 ?linear equation

OpenStudy (saifoo.khan):

Sure!

OpenStudy (saifoo.khan):

what help u want?

OpenStudy (anonymous):

i need too find the y but the equation =x

OpenStudy (saifoo.khan):

Ok, shuffle.

OpenStudy (anonymous):

y=x/11-4/11

OpenStudy (saifoo.khan):

11y = x - 4 \[\Huge y = \frac{1x}{11} - \frac{4}{11}\]

OpenStudy (anonymous):

is that the furthest i can go ?